OpenStudy (anonymous):

Determine the domain of the function. f(x)=sqaureroot 7 +x please help me!

OpenStudy (prettygirl_shynice):

okay ill help

OpenStudy (prettygirl_shynice):

http://openstudy.com/study#/updates/521167d8e4b0450ed75d2e7f

OpenStudy (prettygirl_shynice):

got it

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x)=\sqrt{x+7} }\) Domain, is the (set of) values that you can plug instead for x, into the function. Since square roots can't have a negative inside of them, therefore, must be that \(\color{#000000 }{ \displaystyle x+7\ge 0 }\).

OpenStudy (solomonzelman):

And if, \(\color{#000000 }{ \displaystyle x+7\ge 0 }\) then \(\color{#000000 }{ \displaystyle x\ge {\color{red}{?}} }\)

OpenStudy (prettygirl_shynice):

@SolomonZelman good job